# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ Abstract
Climate science needs more efficient ways to study high-impact, low-probability extreme events, which are rare by definition and costly to simulate in large numbers. Rare event sampling (RES) and ensemble boosting offer a novel strategy to extract more information from those occasional simulated events: small perturbations in advance can turn a moderate event into a severe one, which otherwise might not come for many more simulation-years. But how to choose this “advance split time” (AST) remains a challenge for sudden, transient events like precipitation. In this work, we formulate a concrete optimization problem for the AST and instantiate it on an idealized but physically informative model system: a quasigeostrophic turbulent channel flow advecting a passive tracer, which captures key elements of midlatitude storm track dynamics. Three major questions guide our investigation: (1) Can RES methods, in particular ensemble boosting and trying-early adaptive multilevel splitting, accurately sample extreme events of return periods longer than the simulation time? (2) What is the optimal AST, and how does it depend on the definition of the extreme event, in particular the target location? (3) Can the AST be optimized “online” while running RES?
Our answers are tentatively positive. (1) RES can meaningfully improve tail estimation, using (2) an optimal AST of 1-3 eddy turnover timescales, which varies weakly but detectably with target location. (3) A certain functional that we call the thresholded entropy successfully picks out near-optimal ASTs, eliminating the need for arbitrary thresholds that have thus far hindered RES methods. Our work clarifies aspects of the optimization landscape and can, in our view, guide future research efforts on optimizing and sampling transient extreme events more efficiently in general chaotic systems.
@@ 1.1 Background and motivation
The outsize impact of extreme weather events, and the physically exotic processes that cause them, have driven substantial research interest in the tails of climatological probability distributions. The fundamental challenge is scarcity of data: the historical record is too short to enable robust estimation of extremes rarer than a few times per century, even if the climate were stationary. Different modeling paradigms have developed to confront the issue. The most straightforward is direct numerical simulation (DNS), whereby a climate model is integrated extensively and the extreme events tallied, either as a single long run with stationary forcing (e.g., Yeager et al.,, 2006; O’Gorman and Schneider,, 2009) or as an ensemble with non-stationary forcing (e.g., Thompson et al.,, 2017; John et al.,, 2022). This increases the sample size of extreme events, and reduces the relative error (mean/standard deviation) of an empirical estimate $\hat{p}=\frac{\#\text{ extremes}}{N=\#\text{ total samples}}$, but at a slow rate of $\frac{\sqrt{\mathbb{V}[\hat{p}]}}{\mathbb{E}[\hat{p}]}=\frac{\sqrt{p(1-p)/N}}{p}\sim(Np)^{-1/2}$ for $p\ll 1$. For example, estimating a once-per-century storm ($p=0.01$ year-1) to within 10% relative error would take roughly $N=\frac{1}{0.01}(0.10)^{-2}=10^{4}$ model years. Most of that simulation time is wasted, just waiting for the next event.
@@
Rare event sampling (RES) takes a shortcut by repurposing that time to generate more extremes instead—with the tradeoff of having to account for bias properly. RES stands in contrast to many other strategies which, in one way or another, replace the expensive physical model with a cheaper approximation. Extreme value theory gives principles for parametrically estimating distributions tails (Coles,, 2001), but its asymptotic assumptions are not always justified by the finite datasets available, and it is best suited to model univariate distributions (e.g., average temperature over a region) rather than full spatiotemporal processes like storms, although spatial extreme value modeling is steadily progressing (Huser and Wadsworth,, 2022; Huser et al.,, 2025). Hybrid statistical/physical models aim to parameterize physical processes rather than the final output statistics, and include linear inverse models (Penland and Magorian,, 1993); stochastic weather generators based on analogues or Markov state models (van den Dool,, 1989; Ghil et al.,, 2011; Yiou and Jézéquel,, 2020; Finkel et al.,, 2023; Pons et al.,, 2024); empirical downscaling (Vandal et al.,, 2017; Saha and Ravela,, 2024; Rampal et al.,, 2025); statistical (including machine-learned) emulation (Tebaldi et al.,, 2020; Boulaguiem et al.,, 2022); and generative modeling (Mahesh et al., 2024a, ; Mahesh et al., 2024b, ; Watt and Mansfield,, 2024; Sundar et al.,, 2024; Giorgini et al.,, 2024). Generative models in particular are proliferating at a dizzying pace, and they can indeed generate new samples at low cost, but their ability to represent physics outside their training data—perhaps the most essential requirement for extreme event modeling—is rightly regarded with suspicion.
@@
In light of these options, modelers face a tradeoff between bias (incorrect physics or limited resolution) and variance (erratic statistical estimates due to limited sample size). The methods are not mutually exclusive, with many interesting synergies possible (e.g., as conceptualized in Lucente et al.,, 2022), but RES in particular is our focus here as an under-utilized and under-developed strategy to reduce variance without incurring extra bias.
1.2 Rare event sampling: promise, pitfalls, and optimism
@@
This template must be specialized for the kind of target event. Diffusion Monte Carlo, as applied in Ragone et al., (2018) to season-long hot extremes and Webber et al., (2019) to tropical cyclones, performs the split/kill operation at a chronological sequence of time points, extending the timespan of surviving members while aborting discarded members before they can run to completion—thus, before their $R^{*}$ values can even be measured. This is appropriate when the propensity for a future extreme $R^{*}$ is well-approximated by some property $R(\mathbf{x}(t))$ measurable at the present: for example, if $R^{*}$ is the mean temperature from June to August, $R(\mathbf{x}(t))=$ (running average temperature from June 1 to $t$) is a good splitting criterion (Ragone et al.,, 2018). If $R^{*}$ is peak wind speed over a tropical cyclone’s lifetime, $R(\mathbf{x}(t))=$ (minimum sea-level pressure in the eye) is a good splitting criterion (Webber et al.,, 2019).
@@
But suppose that no good predictor exists. In particular, assume that the severity function $R^{*}$ of a simulation is the maximum over time of a user-defined observable $R(\mathbf{x}(t))$, called the intensity function, and that no better predictor for $R^{*}$ is known besides $R$ itself at the present time. In this case, a better choice of RES algorithm might be adaptive multilevel splitting (AMS; Cérou and Guyader,, 2007), or its more general version “trying-early” AMS (TEAMS), which we previously introduced in Finkel and O’Gorman, (2024)—itself a special case of subset simulation (Au and Beck,, 2001) from engineering—in which every ensemble member runs to completion and produces an actual value of $R^{*}$, not some proxy for it. Descendants are then spawned from the ancestor at some advance split time (AST) $A$ before $R^{*}$ is achieved, to give them enough time to diversify and perhaps exceed their ancestor’s severity, but not so much time to forget their ancestor’s special initial conditions. Fig. 1 illustrates this tradeoff when selecting AST in the context of a simple stochastic system, namely Langevin dynamics (Pavliotis,, 2014) with a logarithmic potential,
$dX(t)$ $=\frac{1}{m}Y(t)\,dt$ (1)
$dY(t)$ $=\Big{[}-V^{\prime}(X(t))-\gamma Y(t)\Big{]}\,dt+\sigma\,dW(t)$ (2)
where (3)
$V(x)$ =$=\begin{cases}\frac{\alpha+1}{\beta}\Big{(}\log(\epsilon)+\frac{(x/\epsilon)^{2}-1}{2}\Big{)}&|x|\leq\epsilon\\ \frac{\alpha+1}{\beta}\log|x|&|x|>\epsilon,\end{cases}$ (4)
[…]
@@
There is no general procedure for selecting AST and other hyperparameters, which impedes the application of RES methods to arbitrary target events and models. We have shown empirically in Finkel and O’Gorman, (2024) the existence of an optimal AST—in the sense of accuracy of long return period estimates—that is roughly approximated by the time until $\frac{3}{8}$ of error saturation. But this result might be highly specific to a number of choices made in Finkel and O’Gorman, (2024) with the Lorenz-96 system, in particular relating to
@@
  • The target variable defining intensity (energy density, $x_{k}^{2}$, with $k=0$, though for Lorenz-96 all sites are statistically equivalent).
@@
Practitioners face a vast menu of choices in all four domains, the first two falling under the purview of domain science and the last two falling under algorithm design. If the choice of target variable changes, it stands to reason that the choice of metric should also change, and so any single prescription of AST (like the $\frac{3}{8}$-saturation time) is unlikely to work for all target variables. Error norms incorporating global information will be less relevant than local norms around the target region, and localized error metrics tend to saturate more slowly.
@@
Our primary goal in this study is to establish a general principle for optimizing AST. To explore itspossible dependencies that don’t exist in Lorenz-96, we upgrade to a 2-layer quasigeostrophic (QG) flow with a passive tracer, whose local concentration is our target variable. The 2-layer QG system is paradigmatic minimal model for baroclinic instability, which Lorenz-96 resembles loosely via its Hopf bifurcation structure (van Kekem and Sterk,, 2018), and the tracer represents one important part of the dynamics governing precipitation, namely advection of water vapor; we leave the extra complexity of condensation and latent heating to future work. This way, our study provides a common jumping-off point for other advection-related extremes such as pollution loading (Neelin et al.,, 2010) and even heat waves (Linz et al.,, 2020). This path up the model hierarchy has been trodden before by Qi and Majda, (2016, 2018), who added passive tracers to Lorenz-96 and a QG model respectively and studied extreme fluctuations in the tracer’s Fourier modes. Also, Gálfi et al., (2017) quantified extreme value statistics—including local and global statistics—of QG wind fields themselves. All these works have inspired and guided this one, but we focus distinctly on the link between short-time perturbation dynamics and long-term climate statistics.
@@
The QG model has enough “space” to explore the effects of all four desiderata listed above on optimal AST. In principle, one can do this with an exhaustive suite of experiments: for every target region (location, size) and every version of stochastic input (e.g., perturbation magnitude and spatial scale) of interest, run TEAMS with a wide range of AST parameters, measure the skill of each AST in matching a reference ground truth distribution, and select the optimal AST. In practice, this exhaustive procedure is not feasible, in part because of the huge number of potential targets, but more fundamentally because TEAMS’ performance is highly subject to randomness. Measuring the effect of any parameter change on the algorithm’s performance requires many repetitions—several dozen at least—to average out the variability inherent in Monte Carlo. Moreover, other hyperparameters exist within TEAMS related to “population management”: the number of initial ensemble members, how many of them to kill and clone at every iteration, and the termination criterion, to name a few. Randomness appears not only as physical forcing, but also in selecting which members to clone, thus interacting tightly with the population hyperparameters. One can think of this as sampling bias, which further blurs the imprint of AST itself on performance.
@@
We suspect, however, that AST is a physically intrinsic concept, not just an algorithmic one. Analogously to Lyapunov exponents, which encode the timescale for small perturbations to double, the AST should encode the timescale for extreme values of some target variable to maximize in variability. This statement is heuristic, and our primary goal here is to propose some concrete definitions for it that, like Lyapunov exponents, are intrinsic to the system and don’t depend on arbitrary algorithmic choices. To achieve this goal of defining AST, and to measure it for a range of target variables (which AST may depend on), we have to take on a secondary goal of developing an efficient measuring stick for AST that is likewise independent of algorithmic pecularities. These are our two major contributions.
@@
The rest of the paper is organized as follows. Sect. 2 details the procedure of generating samples and estimating tail statistics, at a model-agnostic level. Sect. 3 specifies the QG system, its numerical simulation, and its extreme value statistics. Sect. 4 specifies the perturbed-ensemble design at a model-specific level and visualizes some examples of perturbed events. Sect. 6 reports the performance of different AST choices, and visualizes the overall “optimization landscape”. Sect. 7 concludes with an outlook and proposed roadmap for subsequent research—theoretical, algorithmic, and applied.

### v2

@@ Abstract
Climate science needs more efficient ways to study high-impact, low-probability extreme events, which are rare by definition and costly to simulate in large numbers. Rare event sampling (RES) and ensemble boosting use small perturbations to turn moderate events into a severe ones, which otherwise might not come for many more simulation-years, and thus enhance sample size. But the viability of this approach hinges on two open questions: (1) are boosted events representative of the yet-unrealized events? (2) How does this depend on the specific form of perturbation, i.e., timing and structure? Timing in particular is crucial for sudden, transient events like precipitation. In this work, we formulate a concrete optimization problem for the advance split time (AST) hyperparameter, and study it on an idealized but physically informative model system: passive tracer fluctuations in a turbulent channel, which captures key elements of midlatitude storm track dynamics. Three major questions guide our investigation: (1) Can RES methods, in particular “ensemble boosting” equipped with a probability estimator and “trying-early adaptive multilevel splitting”, accurately and efficiently sample extreme events? (2) What is the optimal AST, and how does it depend on the event definition, in particular the target location and surrounding flow conditions? (3) Can the AST be optimized “online” while running RES?
Our answers support RES as a viable method: (1) RES can meaningfully improve tail estimation, using (2) an optimal AST of 1-3 eddy turnover timescales depending on location. (3) A “thresholded entropy” statistic is a good proxy for AST optimality, bypassing the tedious threshold-setting that often hinders RES methods. Our work clarifies aspects of the response function of transient extreme events to perturbations, giving a guide for designing efficient, reliable sampling strategies.
@@ 1.1 Background and motivation
The outsize impact of extreme weather events, and the need to understand the physical processes that cause them, have driven substantial research interest in the tails of climatological probability distributions. The fundamental challenge is scarcity of data: the historical record is too short to enable robust estimation of extremes rarer than a few times per century, even if the climate were stationary. Different modeling paradigms have developed to confront the issue. The most straightforward is direct numerical simulation (DNS), whereby a climate model is integrated extensively and the extreme events tallied, either as a single long run with stationary forcing (e.g., Huang et al., 2016; O’Gorman and Schneider, 2009) or as an ensemble with non-stationary forcing (e.g., Thompson et al., 2017; John et al., 2022). This increases the sample size of extreme events, and reduces the relative error (mean/standard deviation) of an empirical estimate $\hat{p}=\frac{\#\text{ extremes}}{N=\#\text{ total samples}}$, but at a slow rate of $\frac{\sqrt{\mathbb{V}[\hat{p}]}}{\mathbb{E}[\hat{p}]}=\frac{\sqrt{p(1-p)/N}}{p}\sim(Np)^{-1/2}$ for $p\ll 1$ (Zuev, 2015). For example, estimating the probability of a once-per-century storm ($p=0.01$ year-1) to within 10% relative error would take roughly $N=\frac{1}{0.01}(0.1)^{-2}=10^{4}$ model years. Most of that simulation time is wasted, just waiting for the next event.
@@
Rare event sampling (RES) takes a shortcut by repurposing that time to generate more extremes instead, perturbing simulations in a targeted way to favor extreme behavior—with the tradeoff of having to account for bias properly. RES was first developed for nuclear safety assessment (Kahn and Harris, 1951), and has since been generalized for diverse applications including structural reliability engineering (Au and Beck, 2001), molecular dynamics (Zuckerman and Chong, 2017), and more recently climate and weather (e.g., Ragone et al., 2018; Webber et al., 2019; Baars et al., 2021). RES stands in contrast to many other strategies which, in one way or another, replace the expensive physical model with a cheaper approximation. Extreme value theory gives principles for parametrically estimating distributions tails (Coles, 2001), but its asymptotic assumptions are not always justified by the finite datasets available, and it is best suited to model univariate distributions (e.g., average temperature over a region) rather than full spatiotemporal processes like storms, although spatial extreme value modeling is steadily progressing (Huser and Wadsworth, 2022; Huser et al., 2025). Hybrid statistical/physical models aim to parameterize physical processes rather than the final output statistics, and include linear inverse models (Penland and Magorian, 1993); stochastic weather generators based on analogues or Markov state models (van den Dool, 1989; Ghil et al., 2011; Yiou and Jézéquel, 2020; Finkel et al., 2023; Pons et al., 2024); empirical downscaling (Vandal et al., 2017; Saha and Ravela, 2024; Rampal et al., 2025); statistical (including machine-learned) emulation (Tebaldi et al., 2020; Boulaguiem et al., 2022; Mahesh et al., 2024a, b); and generative modeling (Watt and Mansfield, 2024; Sundar et al., 2024; Giorgini et al., 2024). Machine learning models in particular are proliferating at a dizzying pace, and they can indeed generate new samples at low cost, but their ability to represent physics outside their training data—perhaps the most essential requirement for extreme event modeling—is rightly regarded with suspicion.
@@
In light of these options, modelers have several tools to help deal with the tradeoff between bias (incorrect physics or limited resolution) and variance (erratic statistical estimates due to limited sample size). The methods are not mutually exclusive, with many interesting synergies possible (e.g., as conceptualized in Lucente et al., 2022), but RES in particular is our focus here as an under-utilized and under-developed strategy to reduce variance without incurring extra bias.
1.2 Rare event sampling: promise and pitfalls
@@
This template must be specialized for the kind of target event. Diffusion Monte Carlo (DMC), as applied to season-long hot extremes (with a variant called “GKTL” after its inventors; Ragone et al., 2018) and tropical cyclones (with a variant called “QDMC” that applies quantile mapping to intensity values; Webber et al., 2019), performs the split/kill operation at a chronological sequence of time points, extending the timespan of surviving members while aborting discarded members before they can run to completion—thus, before their $R^{*}$ values can even be measured. This is appropriate when the propensity for a future extreme $R^{*}$ is well-approximated by some property $R(\mathbf{x}(t))$ measurable at the present: for example, if $R^{*}$ is the mean temperature from June to August, $R(\mathbf{x}(t))=$ (running average temperature from June 1 to $t$) is a good splitting criterion (Ragone et al., 2018). If $R^{*}$ is peak wind speed over a tropical cyclone’s lifetime, $R(\mathbf{x}(t))=$ (minimum sea-level pressure in the eye) is a good splitting criterion (Webber et al., 2019).
@@
But suppose that no good predictor exists. In particular, assume that the severity function $R^{*}$ of a simulation is the maximum over the event’s timespan of a user-defined observable $R(\mathbf{x}(t))$, such as the accumulated rainfall over a small region between t1t-1 day and $t$, which we generically call the intensity function. Assume further that no better predictor for $R^{*}$ is known besides $R$ itself at the present time. In this case, a better choice of RES algorithm might be adaptive multilevel splitting (AMS; Cérou and Guyader, 2007), or more general versions such as “anticipated AMS” (Rolland, 2022) and “trying-early” AMS (TEAMS), which we previously introduced in Finkel and O’Gorman (2024)—itself a special case of subset simulation (Au and Beck, 2001) from engineering—in which every ensemble member runs to completion and produces an actual value of $R^{*}$, not some proxy for it. Descendants are then spawned from the ancestor at some advance split time (AST) $A$ before $R^{*}$ is achieved, to give them enough time to diversify and perhaps exceed their ancestor’s severity, but not so much time to forget their ancestor’s special initial conditions. Fig. 1 illustrates this tradeoff when selecting AST in the context of a simple stochastic system, namely Langevin dynamics (Pavliotis, 2014) with a logarithmic potential which is specified in Appendix A, but the picture alone conveys the essential phenomenon of an optimal AST. The existence of a nontrivial (i.e., strictly positive) optimum is obvious when looking at isolated events, but its precise value is subtle to quantify when our purpose relates to climatological statistics, i.e., averages over many events.
@@
There is no general procedure for selecting AST and other hyperparameters, which impedes the application of RES methods to arbitrary target events and models. We have shown empirically in Finkel and O’Gorman (2024) the existence of an optimal AST—in the sense of accuracy of long return period estimates—that is roughly approximated by the time until $\frac{3}{8}$ of error saturation. But this result might be specific to the Lorenz-96 system and a number of choices made in Finkel and O’Gorman (2024), in particular relating to
  1. 1.
@@
    The target variable defining intensity (energy density, $x_{k}^{2}$, with site index $k=0$, though for Lorenz-96 all sites are statistically equivalent).
  2. 2.
@@
Practitioners face a vast menu of choices in all four domains, the first two falling under the purview of domain science and the last two falling under algorithm design. If the physical model or the choice of target variable changes, it stands to reason that the choice of metric should also change, and any single prescription of AST (like the $\frac{3}{8}$-saturation time) is unlikely to work for all cases. Indeed in our recent application of TEAMS to extremes of temperature and daily precipitation in a general circulation model, we found that the $\frac{3}{8}$ rule provided some guidance but underestimated the optimal AST for both temperature and precipitation (Finkel and O’Gorman, 2025). Error norms incorporating global information will be less relevant than local norms around the target region, which tend to saturate more slowly (Finkel and O’Gorman, 2025).
@@
Our primary goal in this study is to establish a general principle for optimizing AST. To explore its possible dependencies that don’t exist in Lorenz-96, we upgrade to a 2-layer quasigeostrophic (QG) flow with a passive tracer, whose local concentration is our target variable. The 2-layer QG system is paradigmatic minimal model for baroclinic instability in the atmosphere and ocean, which Lorenz-96 resembles loosely via its Hopf bifurcation structure (van Kekem and Sterk, 2018), and the tracer represents one important part of the dynamics governing precipitation, namely advection of water vapor; we leave the extra complexity of condensation and latent heating to future work. This way, our study provides a common jumping-off point for other advection-related extremes such as pollution loading (Neelin et al., 2010) and temperature extremes (Linz et al., 2020). This path up the model hierarchy has been trodden before by Qi and Majda (2016, 2018), who added passive tracers to Lorenz-96 and a QG model respectively and studied extreme fluctuations in the tracer’s Fourier modes. Also, Gálfi et al. (2017) quantified extreme value statistics—including local and global statistics—of QG wind fields themselves. All these works have inspired and guided this one, but we focus distinctly on the link between short-time perturbation dynamics and long-term climate statistics.
@@
The QG model has enough “space” to explore the effects of all four decision axes listed above on optimal AST. In principle, one can do this with an exhaustive suite of experiments: for every target region (location, size) and every version of stochastic input (e.g., perturbation magnitude and spatial scale) of interest, run TEAMS with a wide range of AST parameters, measure the skill of each AST in matching a reference ground truth distribution, and select the optimal AST. In practice, this exhaustive procedure is not feasible, in part because of the huge number of potential targets, but more fundamentally because TEAMS’ performance is highly subject to randomness. Measuring the effect of any parameter change on the algorithm’s performance requires many repetitions—several dozen at least—to average out the variability inherent in Monte Carlo. Moreover, other hyperparameters related to “population management” exist within TEAMS and other rare event algorithms: the number of initial ensemble members, how many of them to kill and clone at every iteration, and the termination criterion, to name a few. Randomness appears not only as physical forcing, but also in selecting which members to clone, thus interacting tightly with the population hyperparameters. One can think of this as confounding due to sampling bias, which further blurs the imprint of AST itself on performance.
So instead of using TEAMS for our investigation, we turn to a related method of ensemble boosting (Gessner et al., 2021; Fischer et al., 2023). The idea of ensemble boosting is simple: identify some extremes from an initial climatic timeseries, and re-simulate them with perturbed antecedent conditions to generate unrealized but physically plausible (and possibly more extreme) scenarios. By focusing on a limited set of ancestor events to boost, we avoid the additional randomness that occurs in TEAMS as the level is raised and additional ensemble members are stochastically added, which simplifies our investigation. In addition, Bloin-Wibe et al. (2025) has developed an approach to estimate probabilities based on the boosted ensembles, and we have also been developing such an estimator that is introduced below. With the addition of an ability to estimate probabilities, ensemble boosting may now be viewed as an RES algorithm.
@@
We suspect that the optimal AST is closely related to a physically intrinsic quantity that is not particular to a given algorithm. Analogously to Lyapunov exponents, which encode the timescale for small perturbations to double, the optimal AST should encode the timescale for extreme values of some target variable to maximize in variability. This statement is heuristic, and a primary goal here is to propose some quantities that are very close to the optimal AST and that, like Lyapunov exponents, are intrinsic to the system and don’t depend on arbitrary algorithmic choices. We propose and evaluate several candidates including metrics based on entropy and expected improvement.
We have three major contributions. First, we develop a new estimator for low probabilities of extreme fluctuations from boosted ensembles, similar to the estimator of Bloin-Wibe et al. (2025) but distinct in the aggregation step. Our approach includes an optional parametric fit of the response function to perturbations (applicable to both estimators), a simple quadratic regression model that imposes regularity on the resulting severity distribution. Second, we use the two estimators to measure the quality of a range of ASTs across a range of target events (tracer concentration at different target locations), finding evidence for an entropy-based optimality principle. Third, and most importantly from a practical perspective, we demonstrate that both estimators successfully approximate low probabilities when the ensembles are launched from a good AST, which the optimality principle can help to select efficiently. Our goal here is not to demonstrate a performant rare event algorithm—only to elucidate a necessary ingredient (AST) to be optimized in future algorithms—but even when comparing statistical errors at equal cost, we find that our boosted ensembles are already competitive with an equal-cost DNS.
@@
The rest of the paper is organized as follows. Sect. 2 details the procedure of generating samples and estimating tail statistics, at a model-agnostic level, and proposes several candidate indicators of measuring ensemble dispersion that may help select an optimal AST. Sect. 3 specifies the QG system, its numerical simulation, and its extreme value statistics. Sect. 4 specifies the perturbed-ensemble design at a model-specific level. Sect. 5 visualizes some examples of perturbed events, and how the AST selection criteria behave on these examples. Sect. 6 reports the performance of different AST choices, and visualizes the overall “optimization landscape”. Sect. 7 concludes with an outlook and proposed roadmap for subsequent research—theoretical, algorithmic, and applied.